OpenStudy (anonymous):

Without drawing the graph of the equation, determine how many points the parabola has in common with the x-axis and whether its vertex lies above, on, or below the x-axis. y = 3x^2 – 12x + 12 A. 1 point in common; vertex on x-axis B. 2 points in common; vertex below x-axis C. 2 points in common; vertex above x-axis D. no points in common; vertex above x-axis

OpenStudy (anonymous):

use the discriminant to work out how many points are in common with the x-axis

OpenStudy (anonymous):

make y=0

OpenStudy (anonymous):

Make y = 0 and then solve to find the discriment of the parabola, then you can find the points you desire.

OpenStudy (raden):

the discriminant is D = b^2 - 4ac according information above, known a = 3, b= -12, and c = 12 so, D = (-12)^2 - 4(3)(12) = 144 - 144 = 0 if the value of discriminant is 0 then parabola has in common with the x-axis only one

OpenStudy (raden):



OpenStudy (raden):

that vertex for this case , on x-axis

OpenStudy (anonymous):

so its D?

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@RadEn